OpenStudy (anonymous):

find the value of x.... 2x=7+4

OpenStudy (zzr0ck3r):

add the 7+4 on the right side and what do you get?

OpenStudy (anonymous):

11 bt what else must i do

OpenStudy (zzr0ck3r):

ok so what does the equation look like after you do that?

OpenStudy (anonymous):

2x=11....

OpenStudy (zzr0ck3r):

ok so now we want to get the x alone. If we divide both sides by 2 what do we have?

OpenStudy (anonymous):

i dnt know

OpenStudy (zzr0ck3r):

so we have 2x = 11 if we divide both sides by two we have \[\frac{ 2x }{ 2 } = \frac{ 11 }{ 2 }\] note: this is the same thing as

OpenStudy (zzr0ck3r):

\[\frac{ 2 }{ 2 }x = \frac{ 2 }{ 11 }\] what is \[\frac{ 2 }{ 2 }\]?

OpenStudy (anonymous):

watever i already got help with my question... your making this too hard for me ...... but i already got help

OpenStudy (zzr0ck3r):

ok, maybe next time tell me what the problem is. And its not nice to "whatever" someone who is spending there time to help you... good luck in the future. the only question I asked you was what is 2/2. I think you know this.

OpenStudy (zzr0ck3r):

@Shanelle

OpenStudy (zzr0ck3r):

....

OpenStudy (anonymous):

your taking long to help me so i jus said what ever........ dats not any bad ..... 2/2 is 1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

im sorry for my behaviour... but thanks for your help

OpenStudy (zzr0ck3r):

np. so now we have 1*x = 11/2 or x = 11/2